OpenStudy (anonymous):

how to do this problem??? find all zeros of t he function and write the polynomial as a product of linear factors?? f(x)= x^4+25x^2+144

myininaya (myininaya):

let u=x^2 so u^2+25u+144=0 solve for u and then take square root to find x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then what i need to do next?

myininaya (myininaya):

after you take square root of u to find x then you have your answer you can probably simplify

OpenStudy (anonymous):

i have the answer are 4i, -4i,3i, -3i... and what is the linear factors ????

myininaya (myininaya):

f(x)=(x-4i)(x+4i)(x-3i)(x+3i)

OpenStudy (anonymous):

thank you so much:)...

myininaya (myininaya):

np :) i check your answers very good!